# Machine learning applications in archaeological practices: a review

### v1

@@ Abstract
Artificial intelligence and machine learning applications in archaeology have increased significantly in recent years, and these now span all subfields, geographical regions, and time periods from prehistory to modern times. The prevalence and success of these applications have remained largely unexamined, as recent reviews on the use of machine learning in archaeology have only focused only on specific subfields of archaeology. Our review examined an exhaustive corpus of 135 articles published between 1997 and 2022. We observed a significant increase in the number of relevant publications from 2019 onwards, mainly concentrated in a few journals and in majority published in open access format. Automatic structure detection and artefact classification were the most represented tasks in the articles reviewed, followed by taphonomy, archaeological predictive modelling, and architectural classification or reconstruction. From the corpus of articles analysed, clustering and unsupervised methods were underrepresented compared to supervised models. Artificial neural networks and ensemble learning account for two thirds of the total number of models used. However, if machine learning models are gaining in popularity they remain subject to misunderstanding. We observed, in some cases, poorly defined requirements and caveats of the machine learning methods used. Furthermore, the goals and the needs of machine learning applications for archaeological purposes are in some cases unclear or poorly expressed. To address this, we propose here a workflow guide for archaeologists to develop coherent and consistent methodologies adapted to their research questions, project scale and available data. As in many areas of modern life, machine learning is rapidly becoming an important tool in archaeological research and practice, particularly useful for the analyses of large and highly multivariate data, although not without limitations. This review highlights the importance of well-defined and well-reported structured methodologies and collaborative practices to maximise the potential of applications of machine learning methods in archaeological research.
@@ 4.2.6 Taphonomic classification
Many of the articles applied a large number of machine learning methods as well as methods that we do not consider machine learning [such as PLSDA and MDA; e.g. Domínguez‑Rodrigo and Baquedano 2018; Abellan et al. 2022] to compare their performance on the specific task, whilst others [e.g. Byeon et al. 2019; Cifuentes-Alcobendas and Domínguez-Rodrigo 2019] were more restrained. This meant that there was no clear machine learning method or family of methods that was used much more frequently than others, although neural networks were applied in all but one of the articles reviewed (cf. 3.3.8), which is not the case for other methods also commonly used in these articles, such as SVMs or random forests.
@@
Some of these methodological issues have already been highlighted previously [McPherron et al., 2022; but see the reply by Abellan et al. 2022; Moclán and Domínguez-Rodrigo 2023], especially dealing with the unclear methodology applied for the process of “bootstrapping” described in one publication also part of our reviewed dataset [Domínguez‑Rodrigo 2018; though note that this is also present in other publications in our dataset: Domínguez‑Rodrigo and Baquedano 2018; Aramendi et al. 2019; Moclán et al. 2019, 2020]. The lack of specificity on how this process was undertaken is serious, since bootstrapping to create a larger overall dataset and then splitting data for validation and testing [as seems to be very explicitly stated in Domínguez‑Rodrigo and Baquedano 2018; Domínguez‑Rodrigo 2018; Aramendi et al. 2019] constitutes a methodological failure, as all the machine learning models will have the same data for both training and testing datasets, rendering any validation results unusable in the same way providing an exam to a student together with its answer sheet would. Although we do not wish to reiterate the discussion on the use or misuse of the term “bootstrapping” from Dominguez-Rodrigo [2018], it is important to highlight how unclear reporting of methodology or code unavailability can lead other researchers to consider the methods applied unsound. For another related example, both Dominguez-Rodrigo [2018] and Domínguez-Rodrigo and Baquedano [2018] report the use of bootstrapping, but then also describe the use of k-fold cross validation performed seemingly alongside it, both of which are part of the caret R package (Kuhn, 2019), but cannot be used together, as seems to be implied. The trainControl command of the caret package is presumably the one used in both articles, even though no code was made available, and no specific package version was specified [except in an appendix of a citation Kuhn 2019, p.556] to confirm any possible changes in the command’s behaviour in the past that would have made it possible. A more generous reading is that this perhaps only applies to one (C5.0 decision trees) of the six algorithms used, for which k-fold cross validation is specifically mentioned, as can be perhaps surmised from Aramendi et al. [2019], who only mention k-folds cross validation specifically for C5.0. Even so, this raises the question of whether k-folds was performed in addition to holdout testing for all methods [as seems to be the case in Domínguez‑Rodrigo and Baquedano 2018; Domínguez‑Rodrigo 2018], or if the C5.0 implementation alone did not use holdout testing, or was the only one to use both [either option possible in Aramendi et al. 2019]. This also exemplifies the importance of carefully reporting the software versions of the programs used, if not making code available outright.
@@
Other issues may require more technical knowledge to identify, such as those in Moclán et al. [2019], another taphonomic classification study suffering from the main bootstrapping issue discussed above, though only partially, as they evaluated all the models used with both a bootstrapped and non-bootstrapped sample. This study sought to predict the agent that caused a specific bone fracture plane in bone fragments (in which many planes can be present) made experimentally; in this case, between anthropic direct percussion, spotted hyena (C. crocuta) carnivory and wolf (C. lupus) carnivory (Moclán et al., 2019). However, the bone fracture plane entries dataset of fracture planes reported in the article’s Electronic Supplementary Material 1 (ESM 1) contain information on the bone fragment they are from, rather than only about the fracture plane itself (Moclán et al., 2019, app. ESM 1). Thus, if the dataset were randomly split, some data points from the same bone fragment used for model validation may have contained identical bone fracture metadata to the training input, allowing the model to use that metadata to predict the agent of the fracture plane, since all bone fracture planes from the same bone fragment were caused by the same agent (fc. Varoquaux and Cheplygina 2022Varoquaux and Cheplygina 2022). In some cases a single bone fragment had six fracture plane data points defined (and thus six data points available to be shuffled into both testing and training), all containing the same bone fragment metadata for the variables of epiphysis presence, length, and number of planes (Moclán et al., 2019, app. ESM 1). Though no code was made available for this publication to examine how these metadata variables were used during model training and validation, the authors do state that they were indeed used (Moclán et al., 2019, p.4668). In addition, the number of entries in ESM 1 do not match the number of fragments reported for any of the three classes (Moclán et al., 2019, app. ESM 1). Even when counting only unique bone fragments instead of planes (using ESM 1), the numbers, while closer to the numbers reported in the Results section, still do not match for any class (Moclán et al., 2019). Furthermore, for the Hyena class, it is also clear some fragments did not have all their fracture planes added to the ESM 1 dataset [e.g. there are an odd number of entries that state the planes per bone fragment is 2, thus implying there is one plane that was not included; Moclán et al. 2019].
@@ 4.3 Recommendations and good practices
Another major concern is the difficulty of clearly delimiting the extent of machine learning. The definition used here was particularly strict, excluding algorithms that are more traditional statistical methods or mainly deal with dimensionality reduction, even if they were reported in many of the articles reviewed here as indeed machine learning, such as k-means, PAM/k-median (Mircea et al., 2015; Altaweel and Squitieri, 2019; Febriawan et al., 2020; Cacciari and Pocobelli, 2021; Demjan et al., 2022; Bouzid and Barge, 2022; Fernee and Trimmis, 2022; Badawy et al., 2022) as well as PCA, LDA, and QDA (Monna et al., 2020; Ma et al., 2021; Abellan et al., 2022; Anglisano et al., 2022; Badawy et al., 2022). It is not easy to draw a line between what is machine learning and what is not, since it extends across the fields of both statistics and computer science, if not more (Bzdok, 2017; Bzdok et al., 2018). Therefore, we do not wish to suggest our definition is conclusive, but we must also highlight the varied understanding of the concept across authors, which can make an analysis of machine learning applications in a specific field challenging.
@@ 4.4 Trends and perspectives
Regarding automatic structure detection, it seems that, without a doubt, these already well represented tasks in our corpus, will likely find economic opportunities. Private and public companies seek to manage costs and risks associated with new installations and construction, and automatic structure detection can help avoid unnecessary excavations. A private company in Ireland has already tested the ADAF model (ÄŒož et al., 2024) to reduce archaeological intervention time during track construction, by predicting unseen and unreferenced archaeological mounds. Archaeological predictive models might also be regarded as potential tools for construction companies or even public administration to regulate their cultural heritage conservation policy. However, it is crucial to use these tools responsibly, acknowledging that machine learning models are not infallible, nor can they replace manual fieldwork. These tools, if applied commercially, ought to be seen as a way to minimise or at the very least better predict the impact of archaeological material during construction, but never as a way to avoid archaeological intervention. In fact, it might be possible to observe or even prevent the destruction of archaeological material (El-Hajj, 2021), especially with widely-verified and trusted methodologies. Both software developers and archaeologists must clearly define thresholds for feature classification or prediction and provide diverse datasets to ensure comprehensive predictions to avoid overlooking archaeological features.

### v2

@@ Abstract
Artificial intelligence and machine learning applications in archaeology have increased significantly in recent years, and these now span all subfields, geographical regions, and time periods from prehistory to modern times. The prevalence and success of these applications have remained largely unexamined, as recent reviews on the use of machine learning in archaeology have only focused only on specific subfields of archaeology. Our review examined an exhaustive corpus of 135 articles published between 1997 and 2022. We observed a significant increase in the number of relevant publications from 2019 onwards, mainly concentrated in a few journals and mainly published in open access format. Automatic structure detection and artefact classification were the most represented tasks in the articles reviewed, followed by taphonomy, archaeological predictive modelling, and architectural classification or reconstruction. From the corpus of articles analysed, clustering and unsupervised methods were underrepresented compared to supervised models. Artificial neural networks and ensemble learning account for two thirds of the total number of models used. However, if machine learning models are gaining in popularity they remain subject to misunderstanding. We observed, in some cases, poorly defined requirements and caveats of the machine learning methods used. Furthermore, the goals and the needs of machine learning applications for archaeological purposes are in some cases unclear or poorly expressed. To address this, we propose here a workflow guide for archaeologists to develop coherent and consistent methodologies adapted to their research questions, project scale and available data. As in many areas of modern life, machine learning is rapidly becoming an important tool in archaeological research and practice, particularly useful for the analyses of large and highly multivariate data, although not without limitations. This review highlights the importance of well-defined and well-reported structured methodologies and collaborative practices to maximise the potential of applications of machine learning methods in archaeological research.
@@ 4.2.6 Taphonomic classification
Many of the articles applied a large number of machine learning methods as well as methods that we do not consider machine learning [such as PLSDA and MDA; e.g. Domínguez-Rodrigo and Baquedano 2018; Abellán et al. 2022] to compare their performance on the specific task, whilst others [e.g. Byeon et al. 2019; Cifuentes-Alcobendas and Domínguez-Rodrigo 2019] were more restrained. This meant that there was no clear machine learning method or family of methods that was used much more frequently than others, although neural networks were applied in all but one of the articles reviewed (cf. 3.3.8), which is not the case for other methods also commonly used in these articles, such as SVMs or random forests.
@@
Some of these methodological issues have already been highlighted previously [McPherron et al., 2022; but see the reply by Abellán et al. 2022; Moclán and Domínguez-Rodrigo 2023], especially dealing with the unclear methodology applied for the process of “bootstrapping” described in one publication also part of our reviewed dataset [Domínguez-Rodrigo 2018; though note that this is also present in other publications in our dataset: Domínguez-Rodrigo and Baquedano 2018; Aramendi et al. 2019; Moclán et al. 2019, 2020]. The lack of specificity on how this process was undertaken is serious, since bootstrapping to create a larger overall dataset and then splitting data for validation and testing [as seems to be very explicitly stated in Domínguez-Rodrigo and Baquedano 2018; Domínguez-Rodrigo 2018; Aramendi et al. 2019] constitutes a methodological failure, as all the machine learning models will have the same data for both training and testing datasets, rendering any validation results unusable in the same way providing an exam to a student together with its answer sheet would. Although we do not wish to reiterate the discussion on the use or misuse of the term “bootstrapping” from Domínguez-Rodrigo [2018], it is important to highlight how unclear reporting of methodology or code unavailability can lead other researchers to consider the methods applied unsound. For another related example, both Domínguez-Rodrigo [2018] and Domínguez-Rodrigo and Baquedano [2018] report the use of bootstrapping, but then also describe the use of k-fold cross validation performed seemingly alongside it, both of which are part of the caret R package (Kuhn, 2019), but cannot be used together, as seems to be implied. The trainControl command of the caret package is presumably the one used in both articles, even though no code was made available, and no specific package version was specified [except in an appendix of a citation Kuhn 2019, p.556] to confirm any possible changes in the command’s behaviour in the past that would have made it possible. A more generous reading is that this perhaps only applies to one (C5.0 decision trees) of the six algorithms used, for which k-fold cross validation is specifically mentioned, as can be perhaps surmised from Aramendi et al. [2019], who only mention k-folds cross validation specifically for C5.0. Even so, this raises the question of whether k-folds was performed in addition to holdout testing for all methods [as seems to be the case in Domínguez-Rodrigo and Baquedano 2018; Domínguez-Rodrigo 2018], or if the C5.0 implementation alone did not use holdout testing, or was the only one to use both [either option possible in Aramendi et al. 2019]. This also exemplifies the importance of carefully reporting the software versions of the programs used, if not making code available outright.
@@
Other issues may require more technical knowledge to identify, such as those in Moclán et al. [2019], another taphonomic classification study suffering from the main bootstrapping issue discussed above, though only partially, as they evaluated all the models used with both a bootstrapped and non-bootstrapped sample. This study sought to predict the agent that caused a specific bone fracture plane in bone fragments (in which many planes can be present) made experimentally; in this case, between anthropic direct percussion, spotted hyena (C. crocuta) carnivory and wolf (C. lupus) carnivory (Moclán et al., 2019). However, the bone fracture plane entries dataset of fracture planes reported in the article’s Electronic Supplementary Material 1 (ESM 1) contain information on the bone fragment they are from, rather than only about the fracture plane itself (Moclán et al., 2019, app. ESM 1). Thus, if the dataset were randomly split, some data points from the same bone fragment used for model validation may have contained identical bone fracture metadata to the training input, allowing the model to use that metadata to predict the agent of the fracture plane, since all bone fracture planes from the same bone fragment were caused by the same agent (cf. Varoquaux and Cheplygina 2022). In some cases a single bone fragment had six fracture plane data points defined (and thus six data points available to be shuffled into both testing and training), all containing the same bone fragment metadata for the variables of epiphysis presence, length, and number of planes (Moclán et al., 2019, app. ESM 1). Though no code was made available for this publication to examine how these metadata variables were used during model training and validation, the authors do state that they were indeed used (Moclán et al., 2019, p.4668). In addition, the number of entries in ESM 1 do not match the number of fragments reported for any of the three classes (Moclán et al., 2019, app. ESM 1). Even when counting only unique bone fragments instead of planes (using ESM 1), the numbers, while closer to the numbers reported in the Results section, still do not match for any class (Moclán et al., 2019). Furthermore, for the Hyena class, it is also clear some fragments did not have all their fracture planes added to the ESM 1 dataset [e.g. there are an odd number of entries that state the planes per bone fragment is 2, thus implying there is one plane that was not included; Moclán et al. 2019].
@@ 4.3 Recommendations and good practices
Another major concern is the difficulty of clearly delimiting the extent of machine learning. The definition used here was particularly strict, excluding algorithms that are more traditional statistical methods or mainly deal with dimensionality reduction, even if they were reported in many of the articles reviewed here as indeed machine learning, such as k-means, PAM/k-median (Mircea et al., 2015; Altaweel and Squitieri, 2019; Febriawan et al., 2020; Cacciari and Pocobelli, 2021; Demjan et al., 2022; Bouzid and Barge, 2022; Fernee and Trimmis, 2022; Badawy et al., 2022) as well as PCA, LDA, and QDA (Monna et al., 2020; Ma et al., 2021; Abellán et al., 2022; Anglisano et al., 2022; Badawy et al., 2022). It is not easy to draw a line between what is machine learning and what is not, since it extends across the fields of both statistics and computer science, if not more (Bzdok, 2017; Bzdok et al., 2018). Therefore, we do not wish to suggest our definition is conclusive, but we must also highlight the varied understanding of the concept across authors, which can make an analysis of machine learning applications in a specific field challenging.
@@ 4.4 Trends and perspectives
Regarding automatic structure detection, it seems that, without a doubt, these already well represented tasks in our corpus, will likely find economic opportunities. Private and public companies seek to manage costs and risks associated with new installations and construction, and automatic structure detection can help avoid unnecessary excavations. A private company in Ireland has already tested the ADAF model (Čož et al., 2024) to reduce archaeological intervention time during track construction, by predicting unseen and unreferenced archaeological mounds. Archaeological predictive models might also be regarded as potential tools for construction companies or even public administration to regulate their cultural heritage conservation policy. However, it is crucial to use these tools responsibly, acknowledging that machine learning models are not infallible, nor can they replace manual fieldwork. These tools, if applied commercially, ought to be seen as a way to minimise or at the very least better predict the impact of archaeological material during construction, but never as a way to avoid archaeological intervention. In fact, it might be possible to observe or even prevent the destruction of archaeological material (El-Hajj, 2021), especially with widely-verified and trusted methodologies. Both software developers and archaeologists must clearly define thresholds for feature classification or prediction and provide diverse datasets to ensure comprehensive predictions to avoid overlooking archaeological features.